# Extremely luminous optical afterglow of a distant and energetic gamma-ray burst GRB 230204B

### v1

@@ 3.3 Afterglow analysis and correlation
Based on the multi-wavelength follow-up observations campaign ( 1.3 ks to 335 days post burst) of GRB 230204B, we noted that BOOTES robotic and DOT observations have a well-sampled light curve in optical / filters. We fitted flux density optical /-band light curves using a simple power-law function with a temporal decay index of -1.85 0.03 (/DOF = 43.47/52). Furthermore, we combined all the optical data (MASTER, BOOTES, DOT, and a few points from GCN) to get a more well-sampled light curve and it could also be fitted with a simple power-law function with a temporal decay index of -1.84 0.02 (/DOF = 81.13/58). We noted that the optical temporal index is steeper than the typically observed decay slope of -1 (Dainotti et al., 2022). On the other hand, the X-ray afterglow light curve of GRB 230204B has very limited (close to the sensitivity of XRT) data points as Swift followed the burst during the late phase and could be fitted with a temporal decay index of -0.72 0.63 (/DOF = 0.58/1).
The correlation between the average optical decay index and optical luminosity provides crucial insights into the physical mechanisms governing GRB afterglows and their surrounding environments. Oates et al. (2012, 2015) examined the possible correlation between the average optical temporal decay indices and optical luminosity (at 200-sec post burst) for GRBs with optical afterglows observed by Swift/UVOT. They found evidence of anti-correlation between these two parameters (200-sec post burst) as shown in Figure 5. This correlation indicates that optical luminous bursts decay more rapidly than optical faint bursts. We measured optical luminosity and temporal decay index for GRB 230204B and overlaid it onto the luminosity-decay correlation. GRB 230204B exhibits a relatively steep decay rate ( 1.4 in MASTER filter) and high optical luminosity (at the first epoch of afterglow detection). We noted GRB 230204B, highlighted in red in the figure, aligns well with the global correlation observed for long-duration GRBs and lies towards the edge (see Figure 5). Possible physical explanations of this correlation are discussed in section 4.2.
@@ 3.4 Broadband afterglow modeling of GRB 230204B
| Model | Top hat | Gaussian | Gaussian Core |
| 0.003 | 0.004 | 0.005 | |
| 0.006 | 0.006 | 0.007 | |
| – | 4.020 | 4.714 | |
| 55.621 | 55.529 | 55.143 | |
| 2.280 | 2.267 | 2.252 | |
| 1.449 | 1.987 | 2.685 | |
| -5.669 | -5.899 | -6.133 | |
| -0.525 | -0.353 | -0.198 | |
| 1.0 | 1.0 | 1.0 | |
| ln () | -101.310.08 | -107.500.08 | -110.530.08 |
| ln (BF) | 7638.16 | 7631.97 | 7628.94 |
@@
The broadband afterglow modeling involved fitting the observed data across multiple wavelengths, including radio, optical, and X-ray bands. Such broadband modeling helps to understand the physical parameters of the jet, such as its opening angle, energy, and the surrounding medium’s density. The afterglow emission is generally explained by the synchrotron radiation model (Sari et al., 1998; Meszaros & Rees, 1997b; Sari & Piran, 1999). In this model, relativistic electrons are accelerated by the GRB shockwave spiral in the magnetic field, emitting radiation across the electromagnetic spectrum. The synchrotron spectrum is characterized by several break frequencies (self-absorption, peak, and cooling frequencies), which evolve with time and affect the observed light curves at different wavelengths.
@@
For the afterglow modeling of GRB 230204B, we utilized afterglowpy333It currently does not account for certain features, including reverse shock emission, an external wind medium (n r-2), or synchrotron self-absorption. (Ryan et al., 2020)/Redback (Sarin et al., 2024) to explore various jet structure models, including the Tophat, Gaussian, and GaussianCore models. The Bayesian inference framework provided by Redback allowed us to test different priors and achieve robust parameter estimation. Specifically, we employed the PyMultiNest sampler to explore the parameter space and used well-defined priors for key parameters such as the observer angle (), core angle (), initial energy (E0), and interstellar medium density (nISM). Due to the limited number of data points, we have fixed bulk Lorentz factor ( 800), constrained using the -Lorentz factor correlation (Liang et al., 2010). The priors used in the analysis include uniform distributions for the energy, core angle, ambient medium density, and a sine prior for the observer’s viewing angle.
@@
The Top-Hat model provided the best fit to the data (see Figure 6), with log evidence (ln ) of -101.31 0.08 and a Bayes factor (ln BF) of 7638.16 as indicated by the Bayesian evidence comparison. These results strongly suggest that the jet structure for GRB 230204B is best described by a simple Top-Hat profile (see Table 3). The posterior distributions (see corner plot as shown in Figure 11 of the appendix) constrained viewing angle of 0.003 rad and a core angle 0.006 rad (limit as no jet break is observed), suggesting a narrow, highly collimated jet structure. The inferred initial energy is relatively high at 55.62 erg, and the ISM density is constrained to 1.45 cm-3.
@@
The Top-Hat jet model provides a robust description of GRB 230204B’s afterglow, indicating a narrow, highly collimated jet. This is consistent with the high fluence observed in the early afterglow phase. The tightly constrained parameters for and indicate that we are likely observing the jet from a viewing angle close to the core, providing a relatively straightforward interpretation in the context of synchrotron emission from a top-hat jet model. A notable feature in the fit is the very low value of , the magnetic field fraction, which was constrained to = -5.67. This low magnetic field can plausibly arise from shock compression of the ambient interstellar medium (ISM), which supports a weakly magnetized medium (Kumar & Barniol Duran, 2009; Wang et al., 2015). Similar values of have been observed in GRBs such as GRB 190530A (Gupta et al., 2022a), GRB 171205A (Leung et al., 2021), GRB 130427A (Fraija et al., 2016), and a few others (Kumar & Barniol Duran, 2009, 2010; Santana et al., 2014), highlighting the diversity in GRB environments. For GRBs, when , there might be Synchrotron Self-Compton (SSC) contribution to afterglow, particularly at higher frequencies (Joshi & Razzaque, 2021). However, in the case of GRB 230204B, the SSC effects appear negligible within the observed wavelength range (i.e., low-frequencies radio and optical), consistent with findings for similar GRBs (Gupta et al., 2022a; Wang et al., 2024). This is further supported by the temporal decay slopes, afterglow modeling (afterglowpy primarily models synchrotron emission processes), and lack of GeV emission (consistent with synchrotron-dominated regimes). Future studies focusing on a broader sample of bright GRBs with low , combined with well-sampled broadband light curves (from gamma-ray to radio) and afterglow modeling that includes SSC contributions, have the potential to shed light on their physical origins and diverse emission mechanisms.
@@ 4.1 Spectral evolution and Possible Radiation Mechanism
The spectral evolution of the prompt emission reveals complex temporal and spectral variability, which provides crucial insights into the nature of the emission mechanisms and the physical conditions in the GRB jet (Bošnjak & Daigne, 2014). Typically, GRB prompt spectra evolve over time, with the peak energy often shifting to lower energies as the burst progresses. This trend, known as “hard-to-soft” evolution, is commonly observed and is thought to reflect cooling processes within the jet (Norris et al., 1986; Bhat et al., 1994). Alternatively, “tracking” behavior, where follows the intensity variations of the burst, has also been observed in some cases (Ryde & Svensson, 1999), suggesting a possible relationship between the central engine activity and the observed emission (Gupta et al., 2021b).
@@
The radiation mechanisms driving GRBs are fundamentally linked to the spectral index observed during the prompt emission phase and the overall radiative efficiency. The spectral index , which describes the shape of the GRB’s spectrum, provides crucial insights into the underlying physical processes. A harder low energy spectral index (i.e., closer to or exceeding -2/3) typically suggests a dominant photospheric or thermal component, where radiation emerges directly from the photosphere of the relativistic outflow. In contrast, a softer a (less than -2/3) often points to non-thermal processes, such as synchrotron emission, where electrons are accelerated in magnetic fields, producing a broad power-law spectrum (Chattopadhyay et al., 2022; Caballero-García et al., 2023).
@@
To investigate the spectral evolution of GRB 230204B, we conducted a time-resolved spectral analysis, selecting time bins based on a Bayesian algorithm for optimal selection. The top panel of Figure 7 presents the deviation information criterion (DIC) values, comparing the Band function with the CPL for each time-resolved bin. The DIC comparison indicates that the majority of bins—excluding three—are better described by the traditional Band model. In the bottom panel of Figure 7, we illustrate the temporal evolution of key spectral parameters, including , and , as derived from the Band and CPL models for GRB 230204B (see Table LABEL:TRS_Table). These parameters are overlaid on the count-rate light curve of the burst’s prompt emission, revealing that the observed exhibits an intensity-tracking behavior. Harder spectral indices during the first and second episodes might suggest a prominent photospheric component, which often produces harder spectra due to thermal-like emission (Ryde et al., 2010). The softer spectral indices during the later bins of the second, third, and fourth episodes might suggest a synchrotron-like emission becomes more prevalent in later times, indicating the coexistence of multiple emission mechanisms as the burst evolves (Li et al., 2021; Acuner & Ryde, 2018).
@@
Additionally, we explored correlations among the spectral parameters, examining (a) log(Flux) versus log(), (b) log(Flux) versus , and (c) log() versus . Pearson correlation analysis was used to quantify the correlation strengths (Pearson correlation coefficient, (r) and to assess the statistical significance (p-value) for each parameter pair. We noted a strong correlation for log(Flux) versus log() of Band and CPL models with correlation coefficients equal to 0.69 and 0.73, respectively. We also noted a moderate correlation for log(Flux) versus of Band model with a correlation coefficient equal to 0.46, but no correlation for log() versus .
@@
The radiation efficiency of GRBs is another key factor, reflecting the fraction of the jet’s kinetic energy converted into gamma rays. It varies significantly depending on the dominant radiation mechanisms within the jet (Wang et al., 2015). In the photospheric model, radiation is produced as photons escape from the optically thick inner regions of the relativistic outflow. Energy is released at the photosphere where the flow becomes transparent, leading to thermal emission. Photospheric models tend to predict higher radiation efficiencies, sometimes exceeding 50-60%, especially if energy dissipation occurs near the photosphere (Lazzati et al., 2009; Beloborodov, 2010; Pe’er & Ryde, 2011; Lazzati et al., 2013). In contrast, the synchrotron models (arising from accelerated particles in internal shocks or magnetic reconnection sites) often predict lower efficiencies, typically in the range of a few percent to–30%, in such cases much of the energy remains in the kinetic form rather than being radiated away (Kobayashi et al., 1997; Spada et al., 2000; Beniamini & Piran, 2013; Kumar & Zhang, 2015). The hybrid model combines thermal photospheric emission with non-thermal synchrotron radiation from shocks or magnetic reconnections within the jet. This model can accommodate both thermal and non-thermal features in GRB spectra. Hybrid models can achieve efficiencies ranging from a few percent to 50% or higher, as they leverage both thermal and non-thermal emission processes (Zhang & Yan, 2011; Zhang & Pe’er, 2009; Vurm et al., 2011). These discussions suggest that GRBs may involve a complex interplay of thermal and non-thermal processes, with the relative contributions of these mechanisms varying across different bursts and even between different phases of the same burst.
@@
The prompt emission and afterglow observations of GRB 230204B present intriguing insights into its jet structure and radiation mechanisms. Our analysis of the prompt emission reveals a spectral hardening during the initial pulses, with a subsequent softening in later pulses, consistent with a hybrid composition (baryonic + Poynting-flux) of Band + Blackbody components. This spectral evolution suggests that the jet may consist of both photospheric (thermal) and synchrotron (non-thermal) components, with the harder spectrum at the start dominated by a photospheric component and later phases increasingly influenced by synchrotron radiation (Sharma et al., 2020; Gupta et al., 2024c). Additionally, our afterglow modeling indicates a low gamma-ray efficiency of 4.3%, suggesting that most of the jet’s energy remains in kinetic form with minimal conversion into prompt gamma-ray emission. While the low gamma-ray efficiency supports a hybrid model, where both thermal photospheric and non-thermal synchrotron components coexist, it does not entirely rule out a synchrotron-only scenario (Rossi et al., 2002). Such a model, without any photospheric thermal emission, could also account for the low efficiency, provided the energy dissipation mechanism predominantly drives non-thermal radiation. However, the observed spectral evolution, with the initial hard component transitioning to softer emission, is more naturally explained by the presence of a photospheric contribution early on. Furthermore, the low gamma-ray efficiency aligns well with the characteristics of an ISM-like top-hat jet structure in the afterglow phase. In such a scenario, a top-hat jet interacting with a homogeneous interstellar medium produces a relatively uniform energy distribution across the jet, reducing the likelihood of highly efficient gamma-ray emission (Sari et al., 1998; Granot & Sari, 2002; Zhang & Kobayashi, 2005). This jet structure also favors an afterglow powered predominantly by forward shock emission, which is consistent with the observed light curve and its steady decay pattern.
@@
The kinetic energy and of GRB 230204B place it within the population of GRBs characterized by relatively low radiation efficiency. The kinetic energy for GRB 230204B, inferred from afterglow modeling, was found to be significantly higher than the isotropic gamma-ray energy, emphasizing that a substantial fraction of the jet’s energy remained in kinetic form and was not radiated as gamma rays. We noted that GRB 230204B has a higher and values than expected from magnetar central engine (2 1052 erg, Metzger et al. 2011). This supports the rapidly rotating black hole as a possible central engine for this energetic and bright burst (Li et al., 2018; Ror et al., 2024). The distribution of radiation efficiencies across the GRB population, as shown in Figure 8, highlights GRB 230204B as the upper edge with a very low efficiency.

### v2

@@ 3.3 Afterglow analysis and correlation
Based on the multi-wavelength follow-up observations campaign ( 1.3 ks to 335 days post burst) of GRB 230204B, we noted that BOOTES robotic and DOT observations have a well-sampled light curve in optical and filters, respectively. We fitted flux density optical and r-bands light curves using a simple power-law function with a temporal decay index of -1.67 0.15 (/DOF = 20.17/47) and -1.83 0.05 (/DOF = 8.25/2), respectively. Furthermore, we combined all the optical data (MASTER, BOOTES, DOT, and a few points from GCN) to get a more well-sampled light curve, and it could also be fitted with a simple power-law function with a temporal decay index of -1.84 0.02 (/DOF = 81.13/58). We noted that the optical temporal index is steeper than the typically observed decay slope of -1 (2022ApJ...940..169D). On the other hand, the X-ray afterglow light curve of GRB 230204B has very limited (the last XRT data point has a marginal detection with 3.53) data points as Swift followed the burst during the late phase and could be fitted with a temporal decay index of -0.72 0.63 (/DOF = 0.58/1).
The correlation between the average optical decay index and optical luminosity provides crucial insights into the physical mechanisms governing GRB afterglows and their surrounding environments. 2012MNRAS.426L..86O; 2015MNRAS.453.4121O examined the possible correlation between the average optical temporal decay indices and optical luminosity (at 200-sec post-burst) for GRBs with optical afterglows observed by Swift/UVOT. They found evidence of anti-correlation between these two parameters (200-sec post-burst) as shown in Figure 6. This correlation indicates that optical luminous bursts decay more rapidly than optical faint bursts. We measured optical luminosity and temporal decay index for GRB 230204B and overlaid it onto the luminosity-decay correlation. GRB 230204B exhibits a relatively steep decay rate ( 1.74 average decay index) and high optical luminosity (at the first epoch of afterglow detection). We noted GRB 230204B, highlighted in red in the figure, aligns well with the global correlation observed for long-duration GRBs and lies towards the edge (see Figure 6). Possible physical explanations of this correlation are discussed in section 4.2.
@@ 3.4 Broadband afterglow modeling of GRB 230204B
| Model | Prior | Top hat | Gaussian | Gaussian Core |
| [0, /2] | 0.003 | 0.004 | 0.005 | |
| [0.0001, 0.1] | 0.006 | 0.006 | 0.007 | |
| [1, 8] | – | 4.020 | 4.714 | |
| [52, 57] | 55.621 | 55.529 | 55.143 | |
| [2.0, 3.5] | 2.280 | 2.267 | 2.252 | |
| ) | [-3, 4] | 1.449 | 1.987 | 2.685 |
| [-7, 0] | -5.669 | -5.899 | -6.133 | |
| [-5, 0] | -0.525 | -0.353 | -0.198 | |
| 1.0 | 1.0 | 1.0 | ||
| ln | -101.310.08 | -107.500.08 | -110.530.08 | |
| ln BF | 7638.16 | 7631.97 | 7628.94 |
@@
The broadband afterglow modeling involved fitting the observed data across multiple wavelengths, including radio, optical, and X-ray bands. Such broadband modeling helps to understand the physical parameters of the jet, such as its opening angle, energy, and the surrounding medium’s density. The afterglow emission is generally explained by the synchrotron radiation model (1998ApJ...497L..17S; mes97; sari99). In this model, relativistic electrons are accelerated by the GRB shockwave spiral in the magnetic field, emitting radiation across the electromagnetic spectrum. The synchrotron spectrum is characterized by several break frequencies (self-absorption, peak, and cooling frequencies), which evolve with time and affect the observed light curves at different wavelengths.
@@
For the afterglow modeling of GRB 230204B, we performed fitting using Redback software (2024MNRAS.531.1203S) via importing afterglow jet models from afterglowpy333It currently does not account for certain features, including reverse shock emission, an external wind medium (n r-2), or synchrotron self-absorption. (2020ApJ...896..166R) to explore various jet structure models, including the Tophat, Gaussian, and GaussianCore models. The Bayesian inference framework provided by Redback allowed us to test different priors and achieve robust parameter estimation. Specifically, we employed the PyMultiNest sampler to explore the parameter space and used well-defined priors for key parameters such as the observer angle (), core angle (), initial energy (E0), and interstellar medium density (nISM). Due to the limited number of data points, we have fixed the bulk Lorentz factor ( 800), constrained using the -Lorentz factor correlation (2010ApJ...725.2209L). The priors used in the analysis include uniform distributions for the energy, core angle, ambient medium density, and a sine prior for the observer’s viewing angle.
@@
The Top-Hat model provided the best fit to the data (see Figure 7), with log evidence (ln ) of -101.31 0.08 and a Bayes factor (ln BF) of 7638.16 as indicated by the Bayesian evidence comparison. These results strongly suggest that the jet structure for GRB 230204B is best described by a simple Top-Hat profile (see Table 3). The posterior distributions (see corner plot as shown in Figure 12 of the appendix) constrained viewing angle of 0.003 rad and a core angle 0.006 rad (limit as no jet break is observed), suggesting a narrow, highly collimated jet structure. The inferred initial energy is relatively high at 55.62 erg, and the ISM density is constrained to 1.45 cm-3. We also modeled the afterglow of GRB 230204B with VegasAfterglow (2025arXiv250710829W), assuming a top-hat jet in an ISM-like medium, and obtained parameters consistent with those derived using afterglowpy. The best-fit values are: rad, rad, erg, , , , and .
The Top-Hat jet model provides a robust description of GRB 230204B’s afterglow, indicating a narrow, highly collimated jet. This is consistent with the high fluence observed in the early afterglow phase. The tightly constrained parameters for and indicate that we are likely observing the jet from a viewing angle close to the core, providing a relatively straightforward interpretation in the context of synchrotron emission from a top-hat jet model. The last XRT data point at 10 days post-burst shows a deviation from the afterglow model prediction. This observation corresponds to a marginal detection (3.53) with a count rate of ct sec-1 (exposure of 4533.2 sec), close to the sensitivity limit of Swift-XRT. Due to the limited photon statistics, we could not reliably constrain the spectral parameters. Considering the excess flux is intrinsic, it could arise from additional physical processes such as late-time energy injection (2006ApJ...647.1238J), refreshed shocks (2022MNRAS.513.2777K), or wavelength-dependent emission components. However, in the absence of further X-ray or multiwavelength coverage at comparable epochs, the true origin of this excess remains uncertain.
@@
A notable feature in the fit is the very low value of , the magnetic field fraction, which was constrained to = -5.67. Such low magnetic fields can possibly arise from shock compression of a weakly magnetized ambient medium (2009MNRAS.400L..75K; 2015ApJS..219....9W), and similar values have been observed in GRBs such as GRB 190530A (Gupta1), GRB 171205A (2021MNRAS.503.1847L), GRB 130427A (2016ApJ...818..190F) and a few others (2009MNRAS.400L..75K; 2010MNRAS.409..226K; 2014ApJ...785...29S), highlighting the diversity in GRB environments. For GRBs, when , there might be Synchrotron Self-Compton (SSC) contribution to afterglow (2021MNRAS.505.1718J). However, in the case of GRB 230204B, our afterglow modeling using Redback and VegasAfterglow (see below) suggests that the SSC effects appear negligible, consistent with findings for similar GRBs (Gupta1; 2024ApJ...969..146W). This is further supported by the temporal decay slopes, and no significant GeV emission (consistent with synchrotron-dominated regimes). However, we cannot completely rule out some SSC contribution, particularly at early radio epochs, where current limited data may lack the sensitivity to detect weak SSC components. Future studies focusing on a broader sample of bright GRBs with low , combined with well-sampled broadband light curves (from gamma-ray to radio) and afterglow modeling that includes SSC contributions, have the potential to shed light on their physical origins and diverse emission mechanisms.
Furthermore, to investigate alternative scenarios, we modeled the afterglow of GRB 230204B using the VegasAfterglow framework (2025arXiv250710829W), a high-performance and versatile code that self-consistently treats forward and reverse shock dynamics, synchrotron and inverse Compton radiation, and arbitrary circumburst density profiles. Unlike commonly used tools such as afterglowpy, VegasAfterglow supports wind-like media, allowing us to probe environments shaped by massive-star mass loss. We performed Markov Chain Monte Carlo (MCMC) fitting of the broadband afterglow light curves under the assumption of a stellar-wind environment (see Table 4) with several jet structures (top-hat, power-law, and Gaussian). The best-fitting wind model with a top-hat jet and resulting posterior distributions for key physical parameters are shown in Figures 13 and 14 of the appendix, respectively. This modeling yields constraints on the isotropic-equivalent energy, jet opening angle, electron distribution index, and other microphysical parameters, indicating a luminous afterglow produced by a narrowly collimated jet.
To discriminate between ISM and wind-like circumburst environments for GRB 230204B, we performed a statistical comparison using reduced values for both the models. For the ISM top-hat model: , and for the Wind top-hat model: . While both reduced values are greater than 1, the ISM model provides a notably better fit. To quantify the statistical significance of this improvement, we applied an F-test. The F-statistic yields (F-value), and corresponds to a p-value of , indicating that the ISM model is statistically significantly better at the confidence level. This indicates that the improvement is not merely due to random fluctuations but reflects a genuine preference for the ISM scenario.
Beyond the statistical framework, we examined the standard closure relations for model discrimination. Using the electron energy distribution indices obtained from our afterglow modeling: (ISM scenario), (wind scenario), we calculated the predicted temporal and spectral indices. For the slow-cooling relations, the closure relations are:
For the ISM-derived , these relations predict or and temporal slopes or for the () and () spectral regimes, respectively. For the wind-derived , we obtain or , and temporal indices or for the analogous spectral regimes. Notably, both ISM and wind predictions systematically underestimate the observed steep optical decay rate. Consequently, the standard closure relations fail to discriminate the environments, primarily due to the unavailability of well-sampled multiwavelength data. In contrast, our comprehensive afterglow modeling framework provides a more robust and physically complete determination that the ISM scenario is preferred.
@@ 4.1 Spectral evolution and Possible Radiation Mechanism
The spectral evolution of the prompt emission reveals complex temporal and spectral variability, which provides crucial insights into the nature of the emission mechanisms and the physical conditions in the GRB jet (2014A&A...568A..45B). Typically, GRB prompt spectra evolve over time, with the peak energy often shifting to lower energies as the burst progresses. This trend, known as “hard-to-soft” evolution, is commonly observed and is thought to reflect cooling processes within the jet (1986ApJ...301..213N; 1994ApJ...426..604B). Alternatively, “tracking” behavior, where follows the intensity variations of the burst, has also been observed in some cases (1999ApJ...512..693R), suggesting a possible relationship between the central engine activity and the observed emission (2021MNRAS.505.4086G).
@@
The radiation mechanisms driving GRBs are fundamentally linked to the spectral index observed during the prompt emission phase and the overall radiative efficiency. The spectral index , which describes the shape of the GRB’s spectrum, provides crucial insights into the underlying physical processes. A harder low-energy spectral index (i.e., closer to or exceeding -2/3) typically suggests a dominant photospheric or thermal component, where radiation emerges directly from the photosphere of the relativistic outflow. In contrast, a softer (less than -2/3) often points to non-thermal processes, such as synchrotron emission, where electrons are accelerated in magnetic fields, producing a broad power-law spectrum (2022ApJ...936...12C; 2023MNRAS.519.3201C).
@@
To investigate the spectral evolution of GRB 230204B, we conducted a time-resolved spectral analysis, selecting time bins based on a Bayesian algorithm for optimal selection. The top panel of Figure 8 presents the deviation information criterion (DIC) values, comparing the Band function with the CPL for each time-resolved bin. The DIC comparison indicates that the majority of bins—excluding three—are better described by the traditional Band model. In the bottom panel of Figure 8, we illustrate the temporal evolution of key spectral parameters, including , and , as derived from the Band and CPL models for GRB 230204B (see Table LABEL:TRS_Table). These parameters are overlaid on the count-rate light curve of the burst’s prompt emission, revealing that the observed exhibits an intensity-tracking behavior. Harder spectral indices during the first and second episodes might suggest a prominent photospheric component, which often produces harder spectra due to thermal-like emission (2010ApJ...709L.172R). The softer spectral indices during the later bins of the second, third, and fourth episodes might suggest that a synchrotron-like emission becomes more prevalent in later times, indicating the coexistence of multiple emission mechanisms as the burst evolves (Liang2021; 2018MNRAS.475.1708A).
@@
Additionally, we explored correlations among the spectral parameters, examining 1. log(Flux) versus log(), 2. log(Flux) versus , and 3. log() versus . Pearson correlation analysis was used to quantify the correlation strengths (Pearson correlation coefficient, r) and to assess the statistical significance (p-value) for each parameter pair. We noted a strong correlation for log(Flux) versus log() of Band and CPL models with correlation coefficients equal to 0.69 and 0.73, respectively. We also noted a moderate correlation for log(Flux) versus of Band model with a correlation coefficient equal to 0.46, but no correlation for log() versus .
@@
The radiation efficiency of GRBs is another key factor, reflecting the fraction of the jet’s kinetic energy converted into gamma rays. It varies significantly depending on the dominant radiation mechanisms within the jet (2015ApJS..219....9W). In the photospheric model, radiation is produced as photons escape from the optically thick inner regions of the relativistic outflow. Energy is released at the photosphere, where the flow becomes transparent, leading to thermal emission. Photospheric models tend to predict higher radiation efficiencies, sometimes exceeding 50-60%, especially if energy dissipation occurs near the photosphere (2009ApJ...700L..47L; 2010MNRAS.407.1033B; 2011ApJ...732...49P; 2013ApJ...765..103L). In contrast, the synchrotron models (arising from accelerated particles in internal shocks or magnetic reconnection sites) often predict lower efficiencies, typically in the range of a few percent to 30%, in such cases much of the energy remains in the kinetic form rather than being radiated away (1997ApJ...490...92K; 2000ApJ...537..824S; 2013ApJ...769...69B; 2015PhR...561....1K). The hybrid model combines thermal photospheric emission with non-thermal synchrotron radiation from shocks or magnetic reconnections within the jet. This model can accommodate both thermal and non-thermal features in GRB spectra. Hybrid models can achieve efficiencies ranging from a few percent to 50% or higher, as they leverage both thermal and non-thermal emission processes (2011ApJ...726...90Z; 2009ApJ...700L..65Z; 2011ApJ...738...77V). These discussions suggest that GRBs may involve a complex interplay of thermal and non-thermal processes, with the relative contributions of these mechanisms varying across different bursts and even between different phases of the same burst.
@@
The prompt emission and afterglow observations of GRB 230204B present intriguing insights into its jet structure and radiation mechanisms. Our analysis of the prompt emission reveals a spectral hardening during the initial pulses, with a subsequent softening in later pulses, consistent with a hybrid composition (baryonic + Poynting-flux) of Band + Blackbody components. This spectral evolution suggests that the jet may consist of both photospheric (thermal) and synchrotron (non-thermal) components, with the harder spectrum at the start dominated by a photospheric component and later phases increasingly influenced by synchrotron radiation (2020MNRAS.493.5218S; 2024ApJ...972..166G). Additionally, our afterglow modeling indicates a low gamma-ray efficiency of 4.3%, suggesting that most of the jet’s energy remains in kinetic form with minimal conversion into prompt gamma-ray emission. While the low gamma-ray efficiency supports a hybrid model, where both thermal photospheric and non-thermal synchrotron components coexist, it does not entirely rule out a synchrotron-only scenario (2002MNRAS.332..945R). Such a model, without any photospheric thermal emission, could also account for the low efficiency, provided the energy dissipation mechanism predominantly drives non-thermal radiation. However, the observed spectral evolution, with the initial hard component transitioning to softer emission, is more naturally explained by the presence of a photospheric contribution early on. Furthermore, the low gamma-ray efficiency aligns well with the characteristics of an ISM-like top-hat jet structure in the afterglow phase. In such a scenario, a top-hat jet interacting with a homogeneous interstellar medium produces a relatively uniform energy distribution across the jet, reducing the likelihood of highly efficient gamma-ray emission (1998ApJ...497L..17S; 2002ApJ...568..820G; 2005ApJ...628..315Z). This jet structure also favors an afterglow powered predominantly by forward shock emission, which is consistent with the observed light curve and its steady decay pattern.
@@
The kinetic energy and of GRB 230204B place it within the population of GRBs characterized by relatively low radiation efficiency. The kinetic energy for GRB 230204B, inferred from afterglow modeling, was found to be significantly higher than the isotropic gamma-ray energy, emphasizing that a substantial fraction of the jet’s energy remained in kinetic form and was not radiated as gamma rays. We noted that GRB 230204B has a higher and values than expected from magnetar central engine (2 1052 erg, 2011MNRAS.413.2031M). This supports the rapidly rotating black hole as a possible central engine for this energetic and bright burst (2018ApJS..236...26L; 2024ApJ...971..163R). The distribution of radiation efficiencies across the GRB population, as shown in Figure 9, highlights GRB 230204B as the upper edge with a very low efficiency.